OpenStudy (anonymous):

If v,w,x,y, and z are positive consecutive integers and v < w < y < z, then what is the value of z-v/y-v? ( can you explain please) A)3 B)2 C) 4/3 D) 6/5 E) -2

OpenStudy (anonymous):

v = n w = n+1 x = n+2 y = n+3 z = n+4 fraction will be: \[\frac{n+4-n}{n+3-n} = \frac{4}{3}\]

OpenStudy (anonymous):

how do u put in to a fraction :o n+4-n/n+3-n

OpenStudy (anonymous):

in the equaion editor type: \frac{numerator}{denominator}

OpenStudy (anonymous):

equation*

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

thank you very much :)